For each of the following independent cases (A-E), compute the missing values in the table: Total Cost Conversion Cost Direct Materials Direct Manufacturing Manufacturing Labor Case Prime Overhead 2,190 2,120 1,500 2,190 $ 1,090 S 3,630 6,830 11,690 6,980 7,850 3,120 9,350 1,910 3,080 5,820 11,510 20,640 3,440
Answer #1

Total manufacturing cost = Manufacturing overheads +Prime cost

Prime cost = Direct labour and direct material

Conversion cost = Direct labor and manufacturing overheads.

Case

Prime cost

Conversion cost

Direct material

Direct labor

Manufacturing overhead

Total Manufacturing cost

A

$    3,280.00

$           4,720.00

$ 2,190.00

$    1,090.00

$    3,630.00

$   6,910.00

B

$    6,830.00

$         11,690.00

$ 2,120.00

$    4,710.00

$    6,980.00

$ 13,810.00

C

$    4,620.00

$           7,850.00

$ 1,500.00

$    3,120.00

$    4,730.00

$   9,350.00

D

$    2,740.00

$           4,990.00

$     830.00

$    1,910.00

$    3,080.00

$   5,820.00

E

$ 11,510.00

$         20,640.00

$ 3,440.00

$    8,070.00

$ 12,570.00

$ 24,080.00

Working

Case

Prime cost

Conversion cost

Direct material

Direct labor

Manufacturing overhead

Total Manufacturing cost

A

=2190+1090

=1090+3630

2190

1090

3630

=2190+1090+3630

B

6830

11690

2120

=6830-2120

6980

=11690+2120

C

=1500+3120

7850

1500

3120

=7850-3120

9350

D

=830+1910

=1910+3080

=5820-3080-1910

1910

3080

5820

E

11510

20640

3440

=11510-3440

=20640-8070

=12570+8070+3440